Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (08 Jun); Correlator hardware (64 BlB - 1) available. Continued work on data rates.
  • Key dates:
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
  • Staff Travel:
    • MR: 02 Aug - 08 Aug
    • RP: 09 Aug - 21 Aug (URSI)
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
  • Visitors:
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA23 in barn being overhauled.
    • Receivers: EA23 gets outfitted last week in August.
      • Perley: RF Switch testing status.
    • Calibration (Perley): Gain compression/expansion still being looked at.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Software testing predominates (see testing below); expect to get 64th BlB pair replaced this week (b108-b-5); population of StB rack 8 this week.
  • Computing (Robnett):
    • Notes:

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue:
    • 0900-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Quick summary:
    • 07-11 July issue: set-and-remember attenuator settings for IF B and C were swapped; this shouldn't have caused grave issues but somewhere within this, some time variability of the gains was caused; it appears that application of the switched power solves these issues; we're in the process of contacting affected external users and preparing 'fixed' data sets to provide to them. In addition, we're continuing to see some issues with phase jumps on specific antennas (e.g., 07, 13, 22); recently 16 and 18 have been repaired (repair is by replacing 128 MHz clock cables in the affected antennas/IFs). Please continue to report. Please follow guidelines here:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Identified possible issues in the noise related to the 3-level phase rotator.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .Needs close-out testing.
  • Correlator: [Clark/Carlson] Delay rocking across bands. Ongoing discussion of implementation; pending Bryan's return
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • [Vrcic et al]; CM Stability; issues over the weekend/looming vacation have raised priority here.
  • [McMullin et al]; Expanded OSRO Testing for D configuration.
  • [Pokorny et al]; CBE data rate testing: behind schedule; will continue to test 'special' projects to clear to begin.
  • [Sowinski et al]; Executor update testing; some phased array testing (Thursday?).
  • [Brentjens et al]; MCAF update testing (holography improvements).
  • [Goss et al]; Sub-band edge effects testing.
  • [Butler et al]; Executor update testing; planetary.
  • [Carlson et al]; V Stokes follow-up.
  • Subarrays: cleanup.
  • 3-bit sampler testing; automatic gain setting.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • HPC large dataset (500 GB)
  • Executor fshift algorithm (new; increments <100 Hz)
  • Holography mode testing.
  • Phased array/astrometry mode testing.
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing (focus on phase jumps).

Commissioning Groups

    • Adv: Astrometry (Mioduszewski):
      • Notes: Begin testing of UVW precision issues (Brunthaler). Further updates.
  • Calibrator Models
    • Notes: A configuration data being collected (3 SBs left).
  • HPC (Bhatnagar/Owen/Golap):
    • Notes: ~500 GB data set obtained.
  • Imaging (Ott/Golap):
    • Notes:
  • Algorithm (Bhatnagar, Owen):


RSRO A configuration status:

Individual reports (updates)


Demo Science Status


  • Butler: VAO in a nutshell:
    Screen shot 2011-08-02 at 8.32.16 AM.png

Misc Actions

  • Old/Low priority
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • [Rupen]: Development of WIDAR guidelines document

-- JosephMcMullin - 2011-08-02
Topic revision: r1 - 2011-08-02, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback